Two spheres made of same material have radii in the ratio 2:1. if both the spheres are at same temperature, then what is the ratio of heat radiation energy emitted per second by them?
(a) 1: 4
(b) 4: 1
(c) 3: 4
(d) 4: 3

Grade:Upto college level

3 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option b

Regards
Vikas TU
3 years ago

As molarity of HCl is given 0.1 this means concentration i.e. no. of moles per volume is 0.1 for HCl and same will be for [H+]. As each molecule of HCl gives out 1 H+ ion on ionisation.

now let us come to the PH part

PH=log[H+]

=log(0.1). so PH = 1 (very acidic as per ph scale at 298K)

now we know at 298K P[H+] + P[OH-] = 14 ( as Kw = 10 raised to the power -14)

so P[OH-] = 13

so -log[OH-] = 13

and then the answer [OH-] = 10 raised to power of -13(negative 13)

concentration of OH- ion is so less that it can easily be neglected as compaired to H+ ion’s concentration.

Thank you

Hope you are satisfied by the answer.

Vikas TU
3 years ago
Initial rate of cooling is given as:  
 
 
Initial rate of cooling is given as:  
 
 
Q1/Q2 = R1^2 / R2^2 
 
 
 
,.............................................................................